A right circular cone is 5.4cm hight and radius of it`s base is 2cm. It is melted and recast into another right circular cone with radius of base as1.5cm. find the hight of new cone formed

Grade:9

1 Answers

Arun
6 years ago
Fear student
 
Make volume equal for both the cones
Hence
1/3 * pi* 2² * 5.4 = 1/3 * pi* 1.5² * h
4* 5.4/2.25 = h
h = 9.6 cm
 
Regards
Arun (askIITians forum expert)